# Estimating the growth rate of a birth and death process using data from a small sample

### v2

@@ 3.3 Application to Real Data
We apply $\hat{r}_{MSE}$, $\hat{r}_{Inv}$, and $\hat{r}_{MCMC}$ to data from [williams2022life, van2021reconstructing, fabre2022longitudinal, mitchell2022clonal] as saved, e.g., in realCloneData[["cloneTrees"]] of the package cloneRate. These were the same 42 data sets analyzed in [Johnson2023]. For each of these data sets, the sample size $n$ was between $10$ and $109$. The estimates of the growth rate are shown in Figure 17.
Usually, it holds that $\hat{r}_{MSE}\leq\hat{r}_{Inv}\leq\hat{r}_{MCMC}$ for the considered data. For most of the data sets, the three estimators give very similar results, although there are a few exceptions where there are larger differences among the methods. As would be expected from the simulation study, the differences among the estimators decreases for increasing sample size $n.$ Overall, we see that, much like the methods proposed by [Johnson2023], the methods $\hat{r}_{Inv}$ and $\hat{r}_{MSE}$ give results on these real data sets that are comparable to those obtained from Markov chain Monte Carlo methods, for less computational cost.
Figure 17: Application to data from [williams2022life, van2021reconstructing, fabre2022longitudinal, mitchell2022clonal]. On the $x$-axis are the names of the clone and the sample size and on the $y$-axis are the estimators.
@@ 4 Discussion
We provide an estimator for the growth rate of a birth and death processes which outperforms the state-of-the-art methods ([williams2022life, mitchell2022clonal, Johnson2023]) with respect to the mean squared error and the bias, especially for small sample sizes. We also show analytically that for large sample sizes, our estimator outperforms an estimate based on internal branch lengths from [Johnson2023].
@@
Our method only applies to exponentially growing populations. However, tumors do not always follow an exponential growth curve ([zahid2021forecasting]). Nevertheless, as noted in [Johnson2023], as long as the population grows exponentially for a short time, our methods should provide a good estimate of that initial exponential growth rate because the coalescence times of the sampled lineages are usually early in the history of the population. Mathematical results in this direction were established by [cheek2022coalescent]. To test the assumption of an exponentially growing population, one could use skyline methods as introduced by [pybus2000integrated], [drummond2005], and [stadlerskyline]. Skyline methods can reconstruct the population size history under different demographic models.
Our methods also rely on the assumption that all individuals have the same birth and death rates. This is a limitation because in cell populations, sometimes several driver mutations accumulate, leading to fitness differences among the cells ([scott2017somatic]). However, it was noted by [Johnson2023] that site frequency spectrum data from the blood cancer subclones studied in [Johnson2023] agrees well with theoretical predictions for an exponentially growing population in which all cells have the same fitness. Finally, our methods do not take into account any spatial structure in the population and therefore may not be directly applicable to solid tumors.
@@ Acknowledgments
5 Appendix
5.1 Derivation of equation (2.7)
In this section, we calculate $\mathbb{E}\big[L_{n}^{in}|H_{n}^{(1)},...,H_{n}^{(n-1)}\big].$ Recall from (2.6) that
$E\big[L_{n}^{in}|H_{n}^{(1)},...,H_{n}^{(n-1)}\big]=\mathbb{E}\bigg[\bigg(\max_{1\leq i\leq n-1}H_{i,n}-H_{1,n}\bigg)+\sum_{i=1}^{n-2}(H_{i,n}-H_{i+1,n})^{+}\Big|H_{n}^{(1)},...,H_{n}^{(n-1)}\bigg].$
Recall that $H_{i,n}=T-\frac{1}{r}(\log(Q_{n})+U_{i,n})$, and the random variables $U_{1,n},\dots,U_{n-1,n}$ are i.i.d. The random variables $H_{1,n},\dots,H_{n-1,n}$ are not i.i.d. because they all depend on $Q_{n}$, but these random variables are exchangeable. Therefore, it holds that $\mathbb{P}(H_{i,n}=H^{(j)}_{n})=\frac{1}{n-1}$ for all $i,j\in\{1,...,n-1\}$. It follows that
$\mathbb{E}\big[H_{i,n}|H_{n}^{(1)},...,H_{n}^{(n-1)}\big]=\frac{1}{n-1}\sum_{i=1}^{n-1}H_{n}^{(i)}=\frac{1}{n-1}\sum_{i=1}^{n-1}H_{i,n}.$
We also calculate
$\mathbb{E}\bigg[\sum_{i=1}^{n-2}(H_{i,n}-H_{i+1,n})^{+}|H_{n}^{(1)},\dots,H_{n}^{(n-1)}\bigg]$ $=\sum_{i=1}^{n-2}\frac{1}{(n-1)(n-2)}\sum_{v=1}^{n-1}\sum_{u=1}^{n-1}(H_{n}^{(u)}-H_{n}^{(v)})^{+}$
$=\frac{1}{n-1}\sum_{v=1}^{n-1}\sum_{u=1}^{n-1}(H_{n}^{(u)}-H_{n}^{(v)})^{+}$
$=\frac{1}{n-1}\sum_{i=1}^{n}\sum_{j=1}^{n}(H_{i,n}-H_{j,n})^{+}.$
Combining these results, we get
$\mathbb{E}\big[L_{n}^{in}|H_{n}^{(1)},\dots,H_{n}^{(n-1)}\big]=H_{n}^{(n-1)}-\frac{1}{n-1}\sum_{i=1}^{n-1}H_{i,n}+\frac{1}{n-1}\sum_{i=1}^{n-1}\sum_{j=1}^{n-1}(H_{i,n}-H_{j,n})^{+},$
which matches (2.7).
5.2 Calculation of $c_{Inv}(n)$
We show here how to obtain the value of $c_{Inv}(n)$. We have
$c_{Inv}(n)=\mathbb{E}\bigg[\frac{1}{S_{n}}\bigg]=\mathbb{E}[(U_{1,n}-U_{2,n})^{+}].$
Therefore,
$c_{Inv}(n)$ $=\int_{-\infty}^{\infty}\mathbb{P}(U_{i,n}\leq x\leq U_{j,n})\>dx$
$=\int_{0}^{\infty}\int_{-\infty}^{\infty}\mathbb{P}(U_{i,n}\leq x\leq U_{j,n}|Q_{n}=q)f_{Q_{n}(q)}\>dx\>dq$
$=\int_{0}^{\infty}\int_{-\infty}^{\infty}\mathbb{P}(U_{i,n}\leq x|Q_{n}=q)\mathbb{P}(x\leq U_{j,n}|Q_{n}=q)f_{Q_{n}(q)}\>dx\>dq$
Recalling (2.3), we have, for $x\geq-\log q$,
$\mathbb{P}(U_{j,n}\geq x|Q_{n}=q)=\int_{x}^{\infty}\frac{1+q}{q}\cdot\frac{e^{u}}{(1+e^{u})^{2}}\>du=\frac{1+q}{q}\cdot\frac{1}{1+e^{x}}.$
Therefore,
$c_{Inv}(n)=\int_{0}^{\infty}\int_{-\log q}^{\infty}\left(1-\frac{1+q}{q}\frac{1}{1+e^{x}}\right)\bigg(\frac{1+q}{q}\frac{1}{1+e^{x}}\bigg)\frac{nq^{n-1}}{(1+q)^{n+1}}\>dx\>dq.$
Now making the substitution $y=1/(1+e^{x})$, so that dy/dx=y(1y)dy/dx=-y(1-y), we get
$c_{Inv}(n)=\int_{0}^{\infty}\bigg(\int_{0}^{q/(1+q)}\left(1-\frac{1+q}{q}y\right)\frac{1}{1-y}\>dy\bigg)\frac{nq^{n-2}}{(1+q)^{n}}\>dq.$
To evaluate the inner integral, we write 1/(1y)1/(1-y) as an infinite series to get
$\int_{0}^{q/(1+q)}\left(1-\frac{1+q}{q}y\right)\frac{1}{1-y}\>dy$ $=\int_{0}^{q/(1+q)}\bigg(\sum_{j=0}^{\infty}y^{j}-\frac{1+q}{q}\sum_{j=0}^{\infty}y^{j+1}\bigg)\>dy$
$=\sum_{j=0}^{\infty}\bigg(\frac{q}{1+q}\bigg)^{j+1}\frac{1}{j+1}-\sum_{j=0}^{\infty}\bigg(\frac{q}{1+q}\bigg)^{j+1}\frac{1}{j+2}$
$=\sum_{j=0}^{\infty}\bigg(\frac{1}{j+1}-\frac{1}{j+2}\bigg)\bigg(\frac{q}{1+q}\bigg)^{j+1}.$
It follows that
$c_{Inv}(n)$ $=\sum_{j=0}^{\infty}\left(\frac{1}{j+1}-\frac{1}{j+2}\right)\int_{0}^{\infty}\frac{nq^{n+j-1}}{(1+q)^{n+j+1}}\>dq$
$=\sum_{j=0}^{\infty}\left(\frac{1}{j+1}-\frac{1}{j+2}\right)\frac{n}{n+j}$
$=\frac{n}{n-1}\sum_{j=0}^{\infty}\bigg(\frac{1}{1+j}-\frac{1}{n+j}\bigg)-\frac{n}{n-2}\sum_{j=0}^{\infty}\bigg(\frac{1}{2+j}-\frac{1}{n+j}\bigg)$
$=\frac{n}{n-1}\sum_{k=1}^{n-1}\frac{1}{k}-\frac{n}{n-2}\sum_{k=2}^{n-1}\frac{1}{k}$
$=\frac{n}{n-2}\bigg(1-\frac{1}{n-1}\sum_{k=1}^{n-1}\frac{1}{k}\bigg).$
5.3 Extended Information about the real data
We show the distribution of $T$ and $n$ in the data from [williams2022life], [van2021reconstructing, fabre2022longitudinal, mitchell2022clonal] as available in the R package cloneRate.
Figure 18: Distribution of $n$ for the real data from [williams2022life], [van2021reconstructing, fabre2022longitudinal, mitchell2022clonal].
Figure 19: Distribution of $T$ for the real data from [williams2022life], [van2021reconstructing, fabre2022longitudinal, mitchell2022clonal].
5.4 Values of the constants
Table 1 displays the values of the constants $c_{Inv}(n)$$c_{MSE}(n),$ and $c_{Bias}(n)$ to two decimal places for certain values of $n$. The table also gives the reciprocals of the 0.025 and 0.975 quantiles of the distribution of $S_{n}$, which are used in calculating the confidence intervals. The code to calculate these quantiles can be found on our GitHub website.
$n$ $c_{Inv}(n)$ $c_{MSE}(n)$ $c_{Bias}(n)$ $1/q_{0.025}$ $1/q_{0.975}$
5 0.80 0.355 0.59 1.73 0.21
6 0.82 0.49 0.66 1.62 0.27
7 0.83 0.58 0.71 1.53 0.32
8 0.84 0.63 0.74 1.49 0.36
9 0.85 0.67 0.76 1.46 0.40
10 0.86 0.70 0.78 1.43 0.44
11 0.86 0.72 0.80 1.40 0.45
12 0.87 0.74 0.81 1.38 0.47
13 0.88 0.76 0.82 1.36 0.49
14 0.88 0.77 0.83 1.35 0.51
15 0.89 0.79 0.84 1.33 0.52
16 0.89 0.80 0.84 1.32 0.53
17 0.89 0.81 0.85 1.31 0.55
18 0.90 0.82 0.86 1.30 0.56
19 0.90 0.82 0.86 1.29 0.57
20 0.90 0.83 0.87 1.28 0.58
30 0.93 0.87 0.92 1.22 0.65
40 0.94 0.90 0.93 1.20 0.71
50 0.95 0.92 0.93 1.18 0.74
60 0.95 0.93 0.95 1.16 0.76
70 0.96 0.94 0.96 1.15 0.78
80 0.96 0.94 0.96 1.14 0.79
90 0.96 0.95 0.96 1.13 0.80
100 0.97 0.95 0.96 1.13 0.81
Table 1: Values of the constants.
5.5 Influence of the maximal growth rate on the estimator
In Figure 20, we present the RMSE for $n=5,10,20$ and $r=1$ with different values of the maximal growth rate ($4,5,10$), using the R-function birthDeathMCMC of the R package cloneRate. There, we see that performance improves, especially for small sample sizes, drastically when we reduce the maximal growth rate from 10 to 5. We performed 100 simulation runs for each case.
Figure 20: Influence of the value for maxGrowthRate in the function birthDeathMCMC in the R-package cloneRate. Here, MGR stands for maximal growth rate.

### v3

@@ 3.3 Application to Real Data
We apply $\hat{r}_{MSE}$, $\hat{r}_{Inv}$, and $\hat{r}_{MCMC}$ to data from [29, 28, 8, 17] as saved, e.g., in realCloneData[["cloneTrees"]] of the package cloneRate. These were the same 42 data sets analyzed in [12]. For each of these data sets, the sample size $n$ was between $10$ and $109$. The estimates of the growth rate are shown in Figure 17.
Usually, it holds that $\hat{r}_{MSE}\leq\hat{r}_{Inv}\leq\hat{r}_{MCMC}$ for the considered data. For most of the data sets, the three estimators give very similar results, although there are a few exceptions where there are larger differences among the methods. As would be expected from the simulation study, the differences among the estimators decreases for increasing sample size $n.$ Overall, we see that, much like the methods proposed by [12], the methods $\hat{r}_{Inv}$ and $\hat{r}_{MSE}$ give results on these real data sets that are comparable to those obtained from Markov chain Monte Carlo methods, for less computational cost.
Figure 17: Application to data from [29, 28, 8, 17]. On the $x$-axis are the names of the clone and the sample size and on the $y$-axis are the estimators.
@@ 4 Discussion
We provide an estimator for the growth rate of a birth and death processes which outperforms the state-of-the-art methods ([29, 17, 12]) with respect to the mean squared error and the bias, especially for small sample sizes. We also show analytically that for large sample sizes, our estimator outperforms an estimate based on internal branch lengths from [12].
@@
Our method only applies to exponentially growing populations. However, tumors do not always follow an exponential growth curve ([31]). Nevertheless, as noted in [12], as long as the population grows exponentially for a short time, our methods should provide a good estimate of that initial exponential growth rate because the coalescence times of the sampled lineages are usually early in the history of the population. Mathematical results in this direction were established by [3]. To test the assumption of an exponentially growing population, one could use skyline methods as introduced by [20], [5], and [25]. Skyline methods can reconstruct the population size history under different demographic models.
Our methods also rely on the assumption that all individuals have the same birth and death rates. This is a limitation because in cell populations, sometimes several driver mutations accumulate, leading to fitness differences among the cells ([23]). However, it was noted by [12] that site frequency spectrum data from the blood cancer subclones studied in [12] agrees well with theoretical predictions for an exponentially growing population in which all cells have the same fitness. Finally, our methods do not take into account any spatial structure in the population and therefore may not be directly applicable to solid tumors.